OpenStudy (anonymous):

Simplify completely: the quantity 14 times x to the 5th power times y to the 4th power plus 21 times x to the third power times y to the 2nd power all over 7 times x to the third power times y 2x8y5 + 3x6y3 2x2y4 + 3xy 2x2y3 + 3y 2x2y3 + 3y2

OpenStudy (campbell_st):

so the question appear to be \[\frac{14x^5y^4 + 21x^3y^2}{7x^3y} = \frac{14x^5y^4}{7x^3y} + \frac{21x^3y^2}{7x^3y}\] look for the common factors that can be removed.

OpenStudy (anonymous):

i don't get it

OpenStudy (campbell_st):

ok... I'll simplify it further \[\frac{14}{7} \times \frac{x^5}{x^3} \times \frac{y^4}{y^1} + \frac{21}{7} \times \frac{x^3}{x^3} \times \frac{y^2}{y^1}\] the index law for division of the same base is subtract the powers \[\frac{x^a}{x^b} = x^{a - b}\] just simplify the numbers.... 14/7 and 21/ 7 and apply the index law.

OpenStudy (anonymous):

sorry i cant get it

OpenStudy (campbell_st):

ok... so here we go.. try and simplify this \[14\div7 \times x^{5 - 3}\times y^{4 -1} + 21\div7 \times x^{3-3}y^{2-1}\]

OpenStudy (anonymous):

2x^{2}y^{3}+3xy^{1}

OpenStudy (campbell_st):

well almost.... the law for a zero power is \[x^0 = 1\] so the 2nd term does not have an x since \[x^{3 -3} = 1\]

OpenStudy (campbell_st):

hope this has all helped.

OpenStudy (anonymous):

so the answer is wht?

OpenStudy (campbell_st):

ok... so the 1st part is correct \[2x^2y^3 + 3 \times 1\times y\] just simplify for your answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

It's 2x^y^3 + 3y